create t accounts and a partial income statement using the following info:At the beginning of the current season on April 1. the edge of Cullumber Pro Shop showed Cash $3.065, Inventory $4.065 and Co
Answer #1

T-accounts

Cash:-

Date Details Amount($) Date Details Amount($)
1-Apr. Balance 3065 7-Apr. Freight in 90
30-Apr. Account receivable 1379 14-Apr. Account payable 1261
21-Apr. Account payable 784
30-Apr. Balance 2309
4444 4444

Note :- 14 April = 1695-395=1300*3%=39

1300-39=1261

21 April = 938-138=800*2%=16

800-16=784

Inventory account:-

Date Details Amount ($) Date Details Amount ($)
1-apr. Balance 4065 30-apr. Revenue 4065
4065 4065

Common stock:-

Date Details Amount ($) Date Details Amount ($)
30-apr. Balance 7130 1-apr. Balance 7130
7130 7130

Purchase account:-

Date Details Amount ($) Date Details Amount ($)
5-apr. Account payable 1695
12-apr. Account payable 938 30-apr. Balance 2633
2633 2633

Account payable:-

Date Details Amount ($) Date Details Amount ($)
9-apr. Purchase return 395 5-apr. Purchase 1695
14-apr. Purchase discount 39 12-apr purchase 938
14-apr. Cash 1261
17-apr. Purchase return 138
21-apr purchase discount 16
21-apr. Cash 784
2633 2633

Freight account:-

Date Details Amount ($) Date Details Amount ($)
7-apr. Cash 90 7-apr. Balance 90
90 90

Purchase return:-

Date Details Amount ($) Date Details Amount ($)
9-apr Account payable 395
30-apr. Balance 533 17-apr Account payable 138
533 533

Account receivables:-

Date Details Amount ($) Date Details Amount ($)
10-apr. Sales revenue 1514 27-apr. Sales return 90
20-apr. Sales revenue 915 30-apr Cash 1379
30-apr. Balance 960
2429 2429

Sales Revenue :-

Date Details Amount ($) Date Details Amount ($)
10-apr. Account receivable 1514
30-apr. Balance 2429 20-apr. Account receivable 915
2429 2429

Purchase discount:-

Date Details Amount ($) Date Details Amount ($)
14-apr. Account payable 39
30-apr. Balance 55 21-apr. Account payable 16
55 55

Sales return:-

Date Details Amount ($) Date Details Amount ($)
27-apr. Account receivable 90 30-apr. Balance 90